# Language Modeling for the Future of Finance: A Quantitative Survey into Metrics, Tasks, and Data Opportunities

### v1

@@ 7 Potential in Financial Forecasting
Category I papers reveal several areas where forecasting models could be improved. Most studies use general ML metrics, leaving out finance-specific measures that better capture practical utility. Crisis periods are rarely considered, limiting insights into how models behave under stress. Finally, there is a strong U.S. and English-language bias in data, with limited adoption of global financial datasets. Expanding these areas of exploration could strengthen model robustness and support broader real-world applicability.
7.1 Potential in Financial Metrics
As shown in Figure 6, Category I papers mostly rely on ML metrics like accuracy, F1, or MSE (Sawhney et al., 2020b; Wu, 2020; Yangjia et al., 2022). While useful, these do not fully reflect financial performance. Only three financial metrics appear more than once – Sharpe Ratio, Maximum Drawdown, and Cumulative Return (defined in Appendix G).
Finance-specific metrics offer risk-adjusted insights that ML measures overlook. Sharpe Ratio, for example, evaluates return relative to risk (Zou et al., 2022). Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021b).
7.2 Crisis Years for Stress Testing
Figure 7: Distribution of data years used in Category I papers, highlighting publication trends, recession periods, and key financial data milestones.
Figure 7 shows that data usage begins to rise after 1993, when electronic filings became publicly available, with another increase in 2005 following the HTML filing mandate (U.S. Securities and Exchange Commission, ). However, usage dips around 2009 and 2020 – years associated with the global financial crisis and COVID-19 pandemic. These periods are essential for evaluating model robustness, yet they remain underused, likely due to data availability challenges prior to 2015 (Bloomberg L.P., 2015).
Despite these constraints, crisis periods offer valuable testing grounds. Most models are trained on stable conditions, which may not reflect their behavior during volatility. Including data from crises enables stress testing (Investopedia Staff, 2025b) and helps build forecasting models that remain reliable under instability.
7.3 Data Sources
Commonly Used and Overlooked Data Sources
Most forecasting models use common financial sources: stock prices, SEC filings, financial news, earnings calls (e.g., MDRM (Qin & Yang, 2019), StockNet (Xu & Cohen, 2018)). But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Menzio et al., 2024)) or completely absent from NLP research. Public resources like FRED (Federal Reserve Bank of St. Louis, 2024) and Fama-French (Fama & French, 2024) offer rich macroeconomic indicators that remain unexplored. Incorporating these datasets could help capture macroeconomic trends and market dynamics, improving model reliability.
Data Accessibility and Maintenance
Access limitations are also a concern. Legacy datasets like MDRM are outdated, and APIs such as Yahoo Finance’s have been discontinued. Others, like CRSP, require paid access. This limits reproducibility and restricts access to reliable, up-to-date data. There’s a clear need for open, regularly updated datasets applicable for financial forecasting.
8 Discussion & Conclusion
This paper presents a systematic, quantitative analysis of NLP research applied to finance. From an initial pool of 374 papers across major NLP venues (Appendix E), we curated a final set of 221 papers after filtering. We categorized them by primary task (Section 3) and assessed each across 11 qualitative and quantitative dimensions (Table 5). Table 3 summarizes key trends and identifies opportunities where language modeling research can further contribute to finance-related problems.
@@ 8 Discussion & Conclusion
PLM/LLM Adoption General-purpose PLMs and LLMs have largely replaced finance-specific and custom architectures. At the same time, the use of statistical NLP and conventional ML continues to decline (Section 4). Revisiting finance-specific models and exploring alternative methods such as graph-based learning or hybrid statistical models could offer improvements for financial tasks (Section 5).
@@
Data Selection and Bias Most studies use English-language datasets and U.S.-based financial sources. While there is growing diversity in the types of data used (e.g., news, filings, social media), much of it is static, with limited updates or adaptation to changing financial contexts. Developing multilingual and globally representative datasets, along with maintaining and updating existing ones, would support better generalization and long-term applicability of models across financial domains (Section 7.3).
@@
Open Accessibility Code and dataset sharing has become more common, especially after 2021, enhancing transparency and reproducibility. Open access should be a standard in NLP-for-finance research when legally feasible. Maintaining functional and up-to-date repositories is key for long-term reproducibility (Section 5).
Implications for NLP & LM Researchers
NLP research applied to finance has largely centered on a few core tasks – sentiment analysis, information extraction, and question answering – while underexploring others like explainability and privacy-preserving methods (Section 3). The rapid adoption of PLMs and LLMs has made it easier to apply language models, but has also led to reduced methodological variety (Section 4). Graph-based models, statistical techniques, and other alternatives remain underused despite their potential (Section 5). While open-source practices have become more common, researchers should be more careful about maintaining functional and up-to-date repositories to ensure long-term reproducibility and meaningful benchmarking (Section 5). In addition, adopting financial-domain metrics can make evaluations more meaningful and aligned with practical goals (Section 7.1). Including crisis periods in datasets (Section 7.2) and improving data selection practices (Section 7.3) can help ensure robustness and consistency. More diverse datasets, especially beyond English and U.S. markets, are also needed to broaden applicability (Section 7.3).
@@
While NLP methods are increasingly used in financial contexts, several challenges remain. Many models are trained on static or incomplete datasets, leading to biases that may affect predictions (Section 7.3). Evaluations often omit real-world considerations such as market volatility and financial risk (Sections 7.2, 7.1). Incorporating financial benchmarks and testing models under stress conditions could improve their practical value. Continued progress in this direction can better align research outcomes with decision-making needs in financial settings.
Ethics Statement
This study does not assess the merit or quality of individual papers. We do not suggest that any category or method is inherently better than others. Our goal is to map the volume and distribution of research applying NLP methods to finance, and to offer a foundation for further study, without making judgments about the value of specific approaches.
Acknowledgments
We did not receive any specific funding for this work. We thank Arnav Hiray and Michael Galarnyk for their comments.
Appendix A Authorship and Funding
Based on the analysis we conducted, we also made a binary classification of the following categories:
  • Authors affiliated with academia
  • Authors affiliated with Industry
  • Funding from industry
  • Funding from government sources
  • Funding from academic sources
Figure 8: This figure displays the correlation between authors from industrial and academic backgrounds within various types of funding such as academic, industrial and governmental.
As seen in the figure 8, we see that there is no major correlation between funding and authorship. Thus it does not imply that if an author with either industrial or academic is present in the list of authors, there will necessarily be funding from the industry/academia/government.
Category Count
Papers with Academic and Industrial authors 78
Papers with only Academic authors 118
Papers with only Industrial authors 25
Papers with Industrial Funding 49
Papers with Governmental Funding 61
Papers with Academic Funding 35
Table 4: Summary of the count of authors and funding within each category.
Appendix B Various Dimensions surveyed
Exploration dimension Description
Primary task The main task addressed in the paper, such as volatility prediction (Wang et al., 2024).
Sub-Tasks Additional tasks that support the primary objective, like using sentiment analysis to enhance stock price predictions (Jain & Agrawal, 2024).
Methodology Techniques applied, ranging from traditional machine learning to deep learning and large language models. See Section 5.
Code and data availability Whether the paper provides open-source code or data, and whether those artifacts are accessible (e.g., active links).
Contribution What type of contribution the paper has such as a dataset, framework, model, evaluation, and so on.
Comparison/baselines The baselines the paper compares their work to.
Data source The types of data used, including financial reports, social media, news articles, or other financial documents.
Metrics Evaluation criteria, including standard ML metrics (accuracy, precision) and finance-specific metrics (Sharpe ratio).
Data years The time periods covered by the data, such as crisis years or more stable financial periods.
Authorship Indicates what affiliation the authors of the paper have (academic, industrial).
Funding source The origin of the research funding (academic, industrial, or governmental) if any.
Table 5: This table outlines the dimensions examined in our analysis of NLP research focused on the financial domain. Dimensions that led to remarkable findings are discussed in the main paper, while an analysis of authorship and funding is provided in Appendix A.
Appendix C Common Methods used
Methodology Description
Statistical NLP Methods like TF-IDF, Bag-of-Words, and n-grams focused on extracting statistical patterns from financial text.
Embeddings Techniques such as Word2Vec, GloVe, and custom embeddings designed to map financial terms into vector spaces, improving downstream tasks.
Conventional ML Algorithms like Logistic Regression, Support Vector Machines (SVM), and Decision Trees, often used for classification and risk prediction.
Deep Learning (DL) Custom neural network architectures tailored for specific financial tasks such as stock price forecasting.
PLMs Encoder-only models (e.g., BERT) and Encoder-Decoder models (e.g., T5) are used for tasks like document summarization and information extraction.
LLMs Models like GPT-3.5 and LLaMA-2 focus on text generation and understanding complex financial language.
Statistical Modeling Involves correlation analysis, Granger causality, etc., to understand relationships in financial data.
Graphs Uses graph structures to model interactions in financial systems.
Knowledge Graphs Integrating structured knowledge within financial tasks to enhance model performance and reliability.
Table 6: This table categorizes the key methodologies into nine groups.“PLMs” stands for Pretrained Language Models, while “LLMs” stands for Large Language Models.
Appendix D Task Description
The following Table 7 and Table 8 describe the task description with examples.
Primary Task Examples & Insights
Finance-Related (Categories IV)
Explainable AI (XAI) refers to methods that make machine learning outputs more interpretable and less of a "black box." These papers show how small input changes can drastically affect model explanations (Sinha et al., 2021) and highlight the instability of tools like LIME in sensitive applications (Burger et al., 2023). In finance, where decisions must be traceable and justifiable, clear and reliable explanation methods are essential for building trust in tasks like fraud detection and risk assessment.
Privacy-preserving methods protect sensitive data while still allowing model training and analysis. Examples include homomorphic encryption for text similarity (Kim et al., 2022), federated learning for distributed training without raw data exchange (Zhao et al., 2024d), and domain adaptation that shares only model parameters (Xiao et al., 2024b). These methods are especially important in finance, where institutions must work with private data while staying compliant and avoiding breaches.
Numerical reasoning involves working with numbers and structured data to solve problems or develop algorithms. This includes solving math word problems using declarative knowledge (Roy & Roth, 2018), learning numeracy through number embeddings (Duan et al., 2021), and pretraining models for verifying tabular claims (Akhtar et al., 2023). In finance, good numerical reasoning supports tasks like market prediction and risk evaluation by improving how models interpret and reason about numbers.
Financial Applications (Category III)
Sentiment analysis helps assess market mood and forecast financial movements using sources like news, social media, and reports. Studies analyze opinions in reports and news (Rodriguez Inserte et al., 2023), investor sentiment on social media (Guo et al., 2023a), and tone in financial texts (Choe et al., 2023). As markets become more influenced by public opinion, sentiment analysis plays a growing role in trading and analysis.
Information extraction (IE) and relation extraction focus on identifying entities and linking them to events or other entities. Examples include extracting financial events from documents (Zheng et al., 2019), signals in reports (Huang et al., 2023), and patterns in social media (Conforti et al., 2022). These tools help connect companies, individuals, and financial instruments to real-world developments (Liou et al., 2021), supporting real-time analysis and prediction.
Question Answering (QA) enables systems to find specific information in financial documents, reports, and databases. QA models support quick access to relevant facts (Mavi et al., 2023), and recent work focuses on calibrating large language models (LLMs) for finance-specific tasks (Zhao et al., 2024c; Theuma & Shareghi, 2024; Addlesee, 2024). This area emphasizes not only building QA tools but also improving their accuracy for use in high-stakes financial contexts.
Table 7: Overview of key primary tasks in Categories IV (Finance-Related) above the mid line and Category III (Financial Applications) below. Each entry summarizes the task’s role and relevance in NLP research applied to finance, with representative examples and practical insights.
Primary Task Examples & Insights
Financial Resources (Category II)
Dataset and Resource Construction involves building structured, labeled financial datasets for model development and evaluation. These papers focus on creating large-scale annotated resources (Chen et al., 2020a), including datasets like Tab-CQA (Liu et al., 2023) and ConvFinQA (Chen et al., 2022c) designed for reasoning over tables and multi-step numerical queries. Other works assemble corpora from Earnings calls (Pardawala et al., 2025), financial reports (Shah et al., 2022), company filings (Zmandar et al., 2022), news (Tang et al., 2023), and government documents (Shah et al., 2023), enabling downstream tasks like entity linking, numerical reasoning, or forecasting.
Fraud detection leverages domain-specific data and models to identify deceptive financial behavior across platforms. Erben & Waldis (2024) identifies financial scams on Instagram using a fine-tuned BERT model deployed via browser extension and REST API. Another approach detects identity fraud through interactive dialogue, employing knowledge graphs and reinforcement learning to expose inconsistencies in claimed personal data (Wang et al., 2019). These systems highlight how NLP architectures can protect users from financial manipulation.
Financial Forecast (Category I)
Stock Price and Volatility Prediction aims to forecast stock movements or market instability using text data. These models use historical data (Sawhney et al., 2021a), news (Ahbali et al., 2022), and event-driven signals (Sawhney et al., 2020a; Wu, 2020) to predict stock trends. Volatility prediction further explores market sensitivity by analyzing unstructured text like press releases or financial articles (Qin & Yang, 2019), helping traders anticipate fluctuations.
Risk Prediction involves identifying the likelihood and impact of adverse financial events, such as defaults or market disruptions. These models analyze unstructured data like earnings call transcripts (Sang & Bao, 2022), regulatory filings, and legal documents to detect early signals of financial risk (Li et al., 2023b). Applications include credit risk estimation, fraud detection, and systemic risk monitoring (Zhang et al., 2024).
Table 8: Descriptions of major primary tasks in Category II (Financial Resources) above the line and Category I (Financial Forecast) below. Each entry summarizes the task’s focus and contribution to NLP research applied to finance, with representative examples.
Appendix E Conference Proceedings
Category Conferences and Workshops
Financial Forecast (Category I) ACL, Australasian Language Technology Association Workshop, CCL, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), Financial Technology and Natural Language Processing (FinNLP) Workshop, ICCL, LREC-COLING, NAACL, NAACL (Industry), SRW
Financial Resources (Category II) ACL, ACL (Industry), CoNLL, EACL, EcoNLP Workshop, EMNLP, EMNLP-IJCNLP, e-Commerce and NLP Workshop, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, GWC, ICON, ICCL, LREC, LREC-COLING, NAACL, NAACL (Industry), NLP4PI Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, SIGHUM Workshop
Financial Applications (Category III) ACL, ATALA, Bridging Human–Computer Interaction and Natural Language Processing Workshop, CCL, CoLM, Computational Approaches to Subjectivity, Sentiment and Social Media Analysis Workshop, DeeLIO Workshop, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), EMNLP-IJCNLP, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, ICON, ICCL, INLG, LREC, LREC-COLING, NAACL, NAACL (Industry), Natural Legal Language Processing Workshop, News Media Content Analysis and Automated Report Generation Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, Safety4ConvAI Workshop, Structured Prediction for Natural Language Processing Workshop, TextGraphs
Finance Related (Category IV) ACL, ACL-IJCNLP, BlackboxNLP Workshop, EMNLP, Financial Technology and Natural Language Processing (FinNLP) Workshop, LREC-COLING, SRW, TrustNLP Workshop
Table 9: Conference and Workshop Categorization
Appendix F Papers per year
Year 2024 2023 2022 2021 2020 2019 2018 2017 2016 2015 2014 2012 2010 2008 2006 2002 2001 1976 1975
Number of Papers 127 53 90 32 57 15 14 21 6 1 5 1 3 3 1 1 1 1 1
Table 10: Number of papers satisfying our filtration criterion per year.
Appendix G Glossary
Metric Definition
Sharpe Ratio Measures how much return an investment gives for each unit of risk, compared to a risk-free asset (Investopedia Staff, 2025a).
Maximum Drawdown The biggest drop from a peak to a low point in a portfolio’s value before it recovers (Investopedia Staff, 2023b).
Cumulative Return The total profit or loss from an investment over time (Investopedia Staff, 2023a).
Table 11: Key Definitions of financial terms used in our paper.

### v2

@@ 4.3 Opportunities in Data Coverage
Regulatory Filings and Earnings Calls
U.S. public companies file a range of documents such as 10-K, 10-Q, 8-K, credit agreements, proxy statements, S-1, S-3, and others are accessible via the SEC’s EDGAR platform, through free API calls and RSS feeds (U.S. SEC, 2024). While similar registries exist (e.g., Canada’s SEDAR, Japan’s EDINET, UK’s Companies House), they are rarely used, reinforcing a U.S.-centric and English-language bias (Section 4.4). Earnings calls, where firms discuss quarterly performance with investors, are another rich data source. Seeking Alpha transcribes  4,500 calls quarterly (Seeking Alpha, 2025); other providers include The Motley Fool (The Motley Fool, 2025), FactSet (FactSet Insight, 2025), Thomson Reuters (Thomson Reuters, 2025), and S&P Global (S&P Global Market Intelligence, 2025), though often behind paywalls. Companies also post transcripts/slides on IR websites. Public datasets include transcriptions and audio, such as S&P 500 calls from 2017 (Qin and Yang, 2019) and 1,213 companies’ data covering 2015-2018 (Li et al., 2020a).
Figure 4: Data year distribution in financial forecasting papers, annotated with major financial events and infrastructure milestones. Highlights underuse of crisis periods despite their importance for model robustness.
Financial News and Social Media
Real-time news from sources like Reuters, Bloomberg, Dow Jones, CNBC, FT, and WSJ is vital for forecasting. The Reuters-21578 dataset (UC Irvine, 2023), with 10,369 articles from 1987, remains a benchmark in text analysis. Larger archives such as FNSPID (Zhou et al., 2024) link 15.7 million articles (1999-2023) to S&P 500 stocks. News is accessible via APIs (e.g., Bloomberg (Bloomberg L.P., 2025a)) and free sources like Yahoo Finance (Yahoo Finance, 2025) and Google News (Google News, 2025). On social media, “FinTwit” and platforms like X and StockTwits are used for sentiment analysis. Key datasets include Sentiment140 (Kaggle Contributor, 2025; Chaudhary, 2020), StockNet (Xu and Cohen, 2018), and Reddit datasets (Kaggle Contributor, 2021; Wang et al., 2024). X’s API offers limited free and full historical access under paid tiers (X (2025), formerly Twitter). StockTwits, with Bullish/Bearish sentiment tagging, offers free and commercial API access (StockTwits, 2025) and is often used in forecasting studies.
Macroeconomic and Monetary Data
Central bank communications (policy statements, minutes, and press releases) offer critical macroeconomic guidance. The U.S. Federal Reserve’s FOMC releases detailed statements and transcripts on interest rates, inflation, and economic outlook (Federal Reserve System, 2025), with archives available on the Fed’s website. Despite their market relevance, such texts remain underused in NLP-based forecasting. Shah et al. (2023) addresses this by introducing an annotated dataset of FOMC communications and proposing a sentence-level classification task (hawkish, dovish, neutral) using fine-tuned transformers. Economic indicators like GDP, CPI, and unemployment figures – issued by the BEA (U.S. BEA, 2025), FRED (Federal Reserve Bank of St. Louis, 2024), Eurostat, IMF, OECD, and the World Bank (The World Bank, 2025) – are mostly numerical but often include press commentary, all freely available via official portals.
Analyst Reports
These are proprietary equity research reports with buy/sell calls and sector outlooks. Data providers such as Thomson Reuters, FactSet, Bloomberg, and S&P I/B/E/S aggregate these reports, typically through subscription services. All-in-one platforms like the Bloomberg Terminal (Bloomberg L.P., 2025b) and S&P Capital IQ (S&P Global, 2025) provide access to equity research, earnings estimates, and corporate event data as part of their premium offerings. There is no comprehensive free source; however, excerpts or key takeaways occasionally surface in news articles or on platforms like Seeking Alpha.
4.4 Data Concerns and Biases
Data Accessibility Concerns
Despite the presence of a large variety of data sources, there are significant limitations in data accessibility. Legacy datasets described in Section 4.3 often become outdated, and APIs for accessing data may be discontinued, as the Yahoo Finance API (EODHD, 2024), or restricted by paywalls. Other essential databases, such as CRSP, require paid access and are unavailable to most individual researchers. These barriers severely limit reproducibility and restrict access to reliable, up-to-date data. As a result, there is a clear need for open, regularly updated financial datasets suitable for forecasting.
Data Language and Country Bias
As shown in the previous sections, financial NLP research is heavily skewed toward English-language data, particularly from U.S. markets. Most forecasting datasets rely on U.S. regulatory filings, earnings calls, and financial news, which is often centered on large-cap firms like those in the S&P 500. This narrow focus limits the applicability of models to non-English and non-U.S. contexts. Figure 6 highlights that while some multilingual datasets exist, they are mostly designed for language modeling rather than financial forecasting. Very few datasets enable forecasting tasks in other languages, with rare exceptions like the bilingual CMIN corpus (Luo et al., 2023). Expanding multilingual, market-diverse datasets is essential for building globally robust forecasting models and mitigating systemic biases in NLP for finance, as also noted by (Jørgensen et al., 2023).
Multimodal Data Scarcity
Despite significant progress in multimodal learning, the domain of NLP in Finance still lacks comprehensive datasets that integrate text, audio, and other critical financial modalities. One of the few existing resources is MAEC, a dataset aligning transcripts with audio from over 3,400 S&P 500 earnings calls, totaling more than 920 hours of speech, specifically designed for financial risk forecasting (Li et al., 2020b). More recently, FinAudio has emerged as the first standardized benchmark focusing on audio-first tasks (ASR on short and long earnings call segments, plus summarization), compiling over 400 hours of financial audio from multiple sources to evaluate Audio-LLM capabilities (Cao et al., 2025). While many existing datasets support exploration of acoustic and textual cues in financial forecasting, truly multimodal resources (those integrating visual data such as stock charts, presentation slides alongside structured financial metrics) remain very limited. Notable exceptions include Mathur et al. (2022b) and Fons et al. (2024), which begin to address this gap. This modality gap highlights the pressing need for richer, tri- or quad-modal datasets to better capture the multifaceted signals that influence market dynamics.
5 Broader Insights for NLP in Finance
5.1 Complementing Foundation Models with Practical Alternatives
Based on our observations from our corpus of research papers, the methodological landscape in NLP for Finance has shifted decisively toward pretrained language models (PLMs) and large language models (LLMs). We group existing approaches into several broad categories, which allows us to track adoption trends and the rise of PLMs, beginning with the use of BERT-based models in finance around 2019 (Araci, 2019a; Yang et al., 2020c), and accelerating rapidly with the emergence of general-purpose LLMs like ChatGPT (Wu et al., 2023; Gao et al., 2023). However, NLP tasks in Finance often pose domain-specific constraints such as interpretability, regulatory transparency, data scarcity, and the need for low-latency systems. These factors make non-PLM methods (such as rule-based classifiers, feature-engineered models, and domain-specific embeddings) not only viable but often preferable in production environments (Shah et al., 2023; Lopez et al., 2021). The rest of this section revisits key methodological alternatives to PLMs, each addressing practical and conceptual needs that remain critical in industrial financial tasks.
Latency and Scalability
Classical NLP pipelines are far more resource-efficient than LLMs. In time-time contexts like news analytics or alert systems, lightweight models (e.g., logistic regression or small CNNs) deliver insights instantly and can scale to massive corpora without requiring GPU infrastructure (Zhai et al., 2019). Systems like RavenPack (Cornez, 2018) and Refinitiv News Analytics (TRNA) (Waeosri, 2020) use rule-based NLP and horizontally scalable architectures to deliver structured news data with sub-second latency, supporting real-time trading and high-throughput analysis.
Interpretability and Compliance
Lexicon- and rule-based methods remain essential in regulated settings. Financial sentiment lexicons (e.g., Loughran-McDonald) (Loughran and McDonald, 2011) and keyword rules are transparent, auditable, and often outperform neural models in compliance or risk-flagging tasks (Hosseini et al., 2018). Lexicon-based surveillance tools, such as those used in communications monitoring (NICE Actimize, 2023) and financial crime compliance (Risk & Compliance, 2021), are widely adopted for their transparency and ease of audit, allowing firms to justify alerts with specific keywords.
Information Extraction and Knowledge Graphs
Named-entity and relation extraction systems power knowledge graphs and schema-driven applications. These systems offer precision, structure, and explainability, which are ideal for fraud detection, Know Your Customer (KYC), and regulatory alignment, where PLMs often lack controllability (Szarvas et al., 2007; Tarnopolski et al., 2019). Financial institutions use rule-based NLP for entity and relation extraction to build and update knowledge graphs in KYC and Anti-Money Laundering (AML) workflows, as seen in solutions from KYCHub (2025) and TigerGraph (2024), enabling automated risk monitoring and explainable link discovery.
Data Efficiency and Domain Adaptability
Feature-based pipelines using TF-IDF, static embeddings, and classical classifiers (SVMs, XGBoost) work well with limited data. They enable explicit domain knowledge integration (such as tagging financial terms or numerical cues) for credit scoring, bankruptcy prediction, and event detection (Li et al., 2018; Khandani et al., 2021; Alanis et al., 2022). Financial firms often use TF-IDF (Simha, 2021) and XGBoost (Kamau, 2024) to build effective models with limited labeled data without requiring costly infrastructure.
Structure and Summarization
Unsupervised techniques like LDA and extractive summarization remain valuable for discovering latent topics or summarizing dense financial documents. These methods are fast, interpretable, and competitive on factual corpora like earnings calls or filings (Blei et al., 2003; Nenkova and McKeown, 2011). Companies like BlackRock, Bank of America (Phoenix Strategy Group, 2025), and American Express (Arora and Radhakrishnan, 2020) apply topic modeling and extractive summarization to earnings calls and filings, using unsupervised or minimally supervised methods for reliable and traceable outputs.
5.2 Implications of Shift Toward General-Purpose Models
General-purpose language models have rapidly become the default choice in NLP applied to Finance, with models like RoBERTa (Liu et al., 2019), GPT-4 (Achiam et al., 2023), and LLaMA-2 (Touvron et al., 2023) adopted far more widely and quickly than domain-specific alternatives. As Figure 6 shows, even earlier versions of FinBERT (Araci, 2019b; Yang et al., 2020b; Liu et al., 2020; Huang et al., 2022) are used more than newer ones, and recent finance-tuned models like FinGPT (Yang et al., 2023) remain rare. This trend reflects how easy it has become to apply strong general models out of the box. For practitioners, this means many models are optimized for benchmarks, not for the realities of finance, where interpretability, latency, or regulation often matter more than raw accuracy.
Figure 5: Trends in code and dataset availability, highlighting the shift toward open-source practices and the growing accessibility of NLP resources for finance.
Figure 6: Timeline of PLM/LLM adoption in NLP research applied to finance, alongside key datasets by task type: C (custom models), QA (question answering), IE (information extraction), OL (other language), and FD (financial documents). For each model, the top date indicates release; the bottom, first usage in the surveyed papers.
@@ 5.2 Implications of Shift Toward General-Purpose Models
Data Selection and Bias Most studies use English-language datasets and U.S. financial sources. While there is growing diversity in the types of data used (e.g., news, filings, social media), much of it is static, with limited updates or adaptation to changing financial contexts. Developing multilingual and globally representative datasets (Section 4.4), along with maintaining and updating existing ones (Section 4.3), would support better generalization and long-term applicability of models across financial domains.
PLM/LLM Adoption General-purpose PLMs and LLMs have largely replaced finance-specific and custom architectures. At the same time, the use of statistical NLP and conventional ML continues to decline. Revisiting finance-specific models and exploring alternative methods such as graph-based learning or hybrid statistical models could offer improvements for financial tasks (Sections 5.1, 5.2).
@@
Open Accessibility Code and dataset sharing has become more common, especially after 2021, enhancing transparency and reproducibility. Open access should be a standard in NLP-for-finance research when legally feasible. Maintaining functional and up-to-date repositories is key for long-term reproducibility (Section 5.3).
This shift is also mirrored in the decline of custom architectures. Prior to 2022, task-specific models were more common: TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020b)); MDRM (Qin and Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021) in volatility prediction (Niu et al., 2023; Mathur et al., 2022a)). Today, these efforts are rarely pursued, which seems like an efficiency gain at first glance. However, as Section 5.1 discusses, real-world financial applications often require design trade-offs, such as transparency for compliance, or fast inference for deployment, and off-the-shelf LLMs do not accommodate them well. The move away from custom models narrows the solution space at a time when financial tasks still demand methodological flexibility.
5.3 Enabling Reproducibility through Open Resources
As the adoption of language models accelerates, the issue of open accessibility in NLP research, especially in finance, becomes more relevant. As shown in Figure 5, open-source practices have become more common in recent years. Before 2021, closed-source code dominated, but the trend has shifted toward sharing code and datasets.
This move enhances transparency, reproducibility, and collaborative progress in NLP research applied to finance (Whited, 2023). Despite this, some papers still include inactive links or empty repositories, limiting reproducibility. At the same time, dataset availability has steadily increased since 2017, pointing to growing awareness around open data. Maintaining functional, up-to-date repositories remains essential to support meaningful benchmarking and model development.
6 Conclusion & Discussion
Drawing on Table 3, our survey shows that sentiment analysis, information extraction, and forecasting have driven NLP in finance, yet critical gaps remain: finance-specific evaluation metrics, stress-testing on crisis data, and truly global, multilingual datasets. These findings underscore both the field’s progress and the work ahead to build resilient, transparent, and inclusive financial NLP solutions. Below, we discuss their implications for researchers and practitioners.
Implications for NLP and Language Modeling Researchers
The application of LMs to financial domains presents challenges that differ from traditional NLP benchmarks. Our findings suggest that domain-specific requirements, such as interpretability, latency, and compliance, make classical or hybrid approaches highly relevant, even amid the LLM wave. The lack of financial evaluation metrics, underuse of crisis-period data, and limited attention to multilingual and global financial corpora signal areas where NLP researchers can contribute meaningfully. Moreover, the scarcity of domain-adapted models points to missed opportunities for more effective domain transfer. Future work should prioritize robustness, transparency, and real-world relevance over benchmark performance alone.
@@ 6 Conclusion & Discussion
For finance professionals and institutions considering NLP solutions, this survey provides guidance on selecting appropriate tools. Off-the-shelf LLMs may appear attractive, but their performance in finance-specific contexts is often limited by lack of customization, interpretability, or latency guarantees. Classical and lexicon-based methods, when tailored to regulatory or operational constraints, can outperform black-box models in compliance or auditing settings. Practitioners should also evaluate model performance using finance-aligned metrics and test under historical stress scenarios. Finally, leveraging multilingual datasets and broadening market coverage can help mitigate geographic and systemic biases in automated decision-making.